e.com/courses/1615625/quizzes/2363774/take Question 5 How many mL are in 0.468 gallons? (1.000 gallons 3.785 O 5.19 x 102 mL 2.57 x 103 mL 1.43 x 103 mL 9.31 x 102 mL 1.77 x 103 mL
Answer #1

1.000 gallon = 3.785 L so

0.468 gallons =

= 0.468 gallons *(3.785 L/1.000 gallon)

= 1.77 L

1L = 1000 mL so

1.77 L = 1.77*1000 = 1.77*103 mL

so option 5 is correct